OpenStudy (jessiegonzales):

MEDAL!!!! 6. Insert <,> B. < C. =

OpenStudy (jessiegonzales):

@nincompoop

OpenStudy (jessiegonzales):

@zepdrix

zepdrix (zepdrix):

Hmm the bottoms are different :O Makes them difficult to compare. Let's find the `least common multiple` of 4 and 7.

OpenStudy (jessiegonzales):

least common multiple is 1?

zepdrix (zepdrix):

no no, thats the `greatest common factor` :)

zepdrix (zepdrix):

if we count by 4's and count by 7's, when do they meet? :)

OpenStudy (jessiegonzales):

28

zepdrix (zepdrix):

4, 8, 12, 16, 20, 24, 28 7, 14, 21, 28 28 ah good!

zepdrix (zepdrix):

So we want a bottom of 28. \[\Large\rm \frac{3}{4}\cdot\color{royalblue}{\frac{7}{7}}\]See how this guy already has a 4 in the bottom? :d so we gotta give him a 7, ya?

OpenStudy (jessiegonzales):

i see

OpenStudy (jessiegonzales):

so i mulitpy 3 and 4? 21/28

zepdrix (zepdrix):

mmm ok looks good :d how bout the other fraction?

OpenStudy (jessiegonzales):

21/49

zepdrix (zepdrix):

\[\Large\rm \frac{4}{7}\cdot\color{royalblue}{\frac{?}{?}}\]

zepdrix (zepdrix):

Hmm this bottom already has a 7. I don't think we wanna give him another 7 in order to get to 28.

OpenStudy (jessiegonzales):

7/3

zepdrix (zepdrix):

So we actually wanna give him a 4, ya?\[\Large\rm \frac{4}{7}\cdot\color{royalblue}{\frac{4}{4}}\]Because the 7 times 4 will give us that 28 that we were looking for.

OpenStudy (jessiegonzales):

ok

zepdrix (zepdrix):

So what do you get for that second fraction? :) Again, multiply top with top, bottom with bottom.

OpenStudy (jessiegonzales):

16/28

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

i think

zepdrix (zepdrix):

Ok good, So instead of comparing \(\Large\rm \frac{3}{4}\) and \(\Large\rm \frac{4}{7}\) we can get a common denominator and compare \(\Large\rm \frac{21}{28}\) and \(\Large\rm \frac{16}{28}\)

OpenStudy (jessiegonzales):

A~

OpenStudy (jessiegonzales):

>

zepdrix (zepdrix):

the first one is bigger? yay good job \c:/

OpenStudy (jessiegonzales):

thankyou <3